a salesman know to sell a product in 3 out of 5 attempts while another salesman in 3 out of 4 attempts when they both try to sell the product find the probability .
1.no sale will be effected
2.both of them will sell the product
1
Expert's answer
2013-05-23T10:01:14-0400
A represents the event that the first salesman is able to sell the product.
A^c represents the event that the first salesman is not able to sell the product.
B represents the event that the second salesman is able to sell the product.
B^c represents the event that the second salesman is not able to sell the product.
It is given that
P(A)=3/5 and P(B)=2/5
Therefore,
P(A^c )=1- 3/5=2/5 and P(B^c )=1-2/5=3/5
The required probabilities are calculated as follows:
P ( no sale will be affected when they both try to sell the product)
= P(A^c B^c )= P(A^c ) P (B^c) (By multiplication theorem)
=2/5*3/5
=6/25
= 0.24